OpenStudy (cherie_magee):

The slope of a line is 1, and the y-intercept is -1. What is the equation of the line written in slope-intercept form? y = x - 1 y = 1 - x y = -x - 1

OpenStudy (harsimran_hs4):

y = mx + c is equation of line where m is slope and c = y-intercept y = 1x -1 = x-1 y = x-1

OpenStudy (shamim):

Equation of a line is y=mx+c